OpenStudy (anonymous):

What is the x-intercept of the line given by the equation 2x +3y = 12 ? A.The x-intercept is 0. B.The x-intercept is 6. C.The x-intercept is 4. D.The x-intercept is 12.

OpenStudy (anonymous):

If you set your y = 0 and solve for x, you will find the x-intercept. 2x + 3(0) = 12 2x = 12 So what is x?

OpenStudy (anonymous):

Any time you want to find an intercept, say your y-intercept, you set the other variable = 0 and solve for the variable you want the intercept of. In this case we set y = 0 to get the x-intercept. 2x + 3y = 12 The (y) here is next to the 3 so we set it = 0 2x + 3(0) = 12 This becomes, 2x = 12 Solving for x, what do we have to do to get x by itself? Do you know?